# Preheating with deep learning

### v1

@@ 1 Introduction
In this study, we apply deep learning techniques to the following example model. We consider the post-inflationary dynamics of standard model Higgs particles abundantly produced by the inflaton, where the Higgs subsequently decays into gauge bosons through $SU(2)\times U(1)$ interactions, leading to the formation of a thermal sector. During the inflaton oscillation period, numerical simulations are often necessary since the production of scalar or gauge bosons after inflation involves non-perturbative dynamics (e.g. preheating)[20, 21, 22, 23, 24, 25]. Since lattice simulations designed to analyze the preheating system have limitations in observing late-time dynamics up to reheating, we attempt to overcome these limitations by employing deep learning techniques.
@@ 3 Post-inflationary weak interactions
The late-time classical reheating process of self-interacting fields or non-abelian fields exhibits a turbulent and self-similar evolution of distribution functions towards equilibrium [15, 36]. The shape of the spectra, along with the self-similar dynamics, can be comprehended within the framework of wave kinetic theory. This suggests the presence of a consistent trend in the evolution of the occupation number, offering a foundation for exploring the occupation number using deep learning methods.
@@ 4 CNN-LSTM time series analysis
where $\tilde{k}\equiv k\tau^{-p}$, $\tau\equiv t/t_{0}$, and $t_{0}$ denotes a fixed point in time. $p$ and $q$ represent the fitting exponents, and $p$ determines how quickly the system approaches equilibrium. Since the right-hand side of Eq. (14) is stationary, it was used as data for fitting, resulting in a good fit when $p\approx 0.16$ and $q\approx 0.63$, as illustrated in Fig 4. The exponents may vary depending on the details of the preheating model and the fitting conditions.
@@ 5 Conclusion
Acknowledgements. J.Y. would like to thank Oleg Lebedev for the useful discussion. This work was performed using HPC resources from the ‘‘Mésocentre’’ computing center of CentraleSupélec, École Normale Supérieure Paris-Saclay and Université Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-scalay.fr/). This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 860881-HIDDeN, and the CNRS-IRP UCMN.

### v2

@@ 1 Introduction
In this study, we apply deep learning techniques to the following example model. We consider the post-inflationary dynamics of standard model Higgs particles abundantly produced by the inflaton, where the Higgs subsequently decays into gauge bosons through $SU(2)\times U(1)$ interactions, leading to the formation of a thermal sector. During the inflaton oscillation period, numerical simulations are often necessary since the production of scalar or gauge bosons after inflation involves non-perturbative dynamics (e.g. preheating)[20, 21, 22, 23, 24, 25, 26, 27, 28]. Since lattice simulations designed to analyze the preheating system have limitations in observing late-time dynamics up to reheating, we attempt to overcome these limitations by employing deep learning techniques.
@@ 3 Post-inflationary weak interactions
The late-time classical reheating process of self-interacting fields or non-abelian fields exhibits a turbulent and self-similar evolution of distribution functions towards equilibrium [15, 38]. The shape of the spectra, along with the self-similar dynamics, can be comprehended within the framework of wave kinetic theory. [^5] See [39, 40] for possible IR- and UV-cascades in the context of non-thermal fixed points in scalar field theories. This suggests the presence of a consistent trend in the evolution of the occupation number, offering a foundation for exploring the occupation number using deep learning methods.
@@ 4 CNN-LSTM time series analysis
where $\tilde{k}\equiv k\tau^{-p}$, $\tau\equiv t/t_{0}$, and $t_{0}$ denotes a fixed point in time. $p$ and $q$ represent the fitting exponents, and $p$ determines how quickly the system approaches equilibrium. Since the right-hand side of Eq. (14) is stationary, it was used as data for fitting, resulting in a good fit when $p\approx 0.16$ and $q\approx 0.63$, as illustrated in Fig. 4. For physical insights on the exponents, see [15, 38]. However, the aim of this work is not to precisely determine the exponents, but to apply deep learning techniques based on the consistent patterns in the distributions, referred to as self-similar evolution or non-thermal fixed points. The exponents may vary depending on the details of the self- or gauge interaction of the Higgs field as well as the fitting conditions. In fact, the Higgs field interacts not only through self- and gauge interactions but also with fermions, and intense fermion reactions are expected during the preheating period [42, 43]. Once the code is updated to include fermions, it will be interesting to directly observe how this affects the self-evolution or non-thermal fixed points in the Higgs distribution.
@@ 5 Conclusion
Acknowledgements. J.Y. would like to thank Oleg Lebedev for useful discussions. This work was performed using HPC resources from the ‘‘Mésocentre’’ computing center of CentraleSupélec, École Normale Supérieure Paris-Saclay and Université Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-scalay.fr/). This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 860881-HIDDeN, and the CNRS-IRP UCMN.